OpenStudy (anonymous):

Solve: 8x − 2(x + 1) = 7x + 8

OpenStudy (anonymous):

x= -10

OpenStudy (anonymous):

thank you

OpenStudy (shaik0124):

8x-2x-2=7x+8 6x-2=7x+8 6x-2+2=7x+8+2 6x=7x+10 6x-6x=7x-6x+10 0=x+10 -10=x+10-10 x=-10

OpenStudy (anonymous):

You're Welcome :)